# Benchmarking is Broken - Don’t Let AI be its Own Judge

### v1

@@ 3 Towards a New Paradigm: From Static Leaderboards to Proctored Exams
The remainder of the paper instantiates some of these principles in a concrete, community-governed prototype, PeerBench, whose architecture and methodological safeguards are presented in Section 4, demonstrating the practicality of our position.
4 PeerBench: A Live, Community-Governed Benchmarking Platform
We introduce PeerBench [^2] A prototype implementation of PeerBench is available at: https://peerbench.ai., a prototype platform implementing the proctored-exam paradigm outlined earlier. It integrates a lightweight, decentralized governance layer and a secure, cryptographically verifiable workflow to assure data quality of tests and a live, contamination-resistant, and reproducible evaluation.
@@ 4.1 Actors
Conceptually, we will simultaneously support different evaluation streams (e.g. maths, coding, translation) for test datasets. Our system involves four types of actors:
Validators author private test suites, query models, grade responses, and peer-review one another. Model providers expose inference endpoints for their models—open or closed—and optionally register for specific evaluation streams. The Coordination server is a neutral service that authenticates uploads, schedules peer review, updates reputations, and publishes the public leaderboard, while storing artefacts in both a database and immutable object storage for auditability. End-users—such as researchers, journalists, regulators, and practitioners—consult the live leaderboard and may apply validator-trust thresholds when interpreting results.
4.2 End-to-End Workflow
Figure 1: Lifecycle of a PeerBench evaluation round.
We will now outline the workflow from the perspective of an individual validator. All published messages shall be signed to guarantee authenticity.
Setup. Validators and model providers register with verifiable credentials (e.g., institutional email) and generate public signing keys. Validators are required to stake collateral, which may be slashed for misconduct. Upon registration, validators receive two initial reputation scores: data quality reputation $\rho_{DQ}$ which is based on the quality of test content they contribute and influences their contribution to model scores, and peer-review reputation $\rho_{PR}$, which is based on the accuracy and reliability of their reviews of others’ test content and influences their weight in peer-review in the future.
  1. S1
    Validator Onboarding. Validators complete identity verification, stake collateral, generate a cryptographic key pair for signing, and are assigned initial values for $\rho_{DQ}$ and $\rho_{PR}$.
Validator Workflow per Round. Validators generate private test prompts, give a preliminary score to models and prepare their data for audit.
  1. R1
    Benchmark test preparation. In each evaluation round, private benchmarking prompts $Q_{i}^{(v)}$ and scoring functions $F_{i}^{(v)}$ are drafted by the validator $v$.
  2. R2
    Hash commit. Validators publish hash commitments of their private test sets $h_{i}=\mathrm{Com}(Q_{i}^{(v)},F_{i}^{(v)})$ to avoid retroactive cherry-picking. The validator further publicly identifies which evaluation stream its test set belongs to.
  3. R3
    Model Evaluation. For each already registered model to be evaluated on this evaluation stream, the validator should make queries with $Q_{i}^{(v)}$ and store the audit log and raw answers $A_{i}^{(v,m)}$. It further outputs commitments $h_{i,v,m}=\mathrm{Com}(A_{i}^{(v,m)})$ to its logs publicly.
  4. R4
    Local scoring. Objective tasks are auto-graded with scoring function $F_{i}^{(v)}$, producing raw scores $s_{i}^{(v,m)}$; subjective tasks which require preference voting will be scored after test publication. A preliminary score can already be published.
  5. R5
    Reveal Subset for Auditing. Validators reveal a random subset $(Q_{j}^{(v)},F_{j}^{(v)})$ of their test suite, selected via a public randomness beacon (e.g., Drand [^3] https://www.drand.love/. They also publish corresponding model answers and open their earlier commitments $h_{j},h_{j,v,m}$ for verification.)
  6. R6
    Online Peer-Review. Each validator retrieves public tests from others and verifies consistency with committed hashes (from R1). Discrepancies are flagged as cheating. Validators re-run the public tests locally as in steps R3-R4. Significant score or model answer mismatch to what the validator reported causes further investigation.
    Finally, the individual score on the public test set and answers are published.
At the End of a Round. Results are finalized, test sets are retired, and reputations are updated.
  1. R7
    Full Test publication. All test data $Q_{i}^{(v)}$, logs, and model responses $A_{i}^{(v,m)}$ are made public for auditing and further research purposes. Consistency with previous hash commits can be checked. Test sets are permanently retired from future evaluations.
  2. R8
    Score calculation. Objective scores are finalized. Subjective scores (e.g., preference-based) are determined via peer voting. All scores are normalized for cross-model comparability.
  3. R9
    Data quality calculation. Validators score each test set $Q_{i}^{(v)}$ according to originality, rule adherence, and agreement. Each set receives a final quality score $DQ(Q_{i}^{(v)})$, weighted by reviewers’ $\rho_{PR}$ reputations. The final score of each model is the average performance on all tests, weighted by the data quality of each test set.
  4. R10
    Reputation update and Slashing. After data quality calculation of each evaluation cycle, the reputation $\rho_{DQ}$ and $\rho_{PR}$ are updated. $\rho_{DQ}$ is updated according to the data quality $DQ(Q_{i}^{(v)})$ of the test set in the new round, and $\rho_{PR}$ is updated according to the proximity of validator $v$’s scoring on other tests to the final data quality of other tests. Suspicious validators with a reputation lower than a certain threshold will be removed and have their collateral slashed.
Timing. A typical round may span approximately one month, allowing for asynchronous participation. To prevent contamination, only models released before the test generation phase are eligible for scoring in the current round.
@@ 4.3 How Our Design Choices Address Common Issues
  • Data contamination & cherry-picking. Validators pre-commit to test sets, which remain private until the round concludes and are never reused. This ensures that training on this data is infeasible if the validator behaves honestly.
    Any collusion or cherry-picking, such as tailoring questions to favored models, can be detected via cross-validator score discrepancies, discouraging misconduct through slashing and reputation penalties. This incentivizes honest behaviour.
  • Cheating on the private dataset. A public source of randomness determines which queries are revealed, preventing validators from anticipating which items will be audited. Hash commitments to both test items and model responses ensure verifiable consistency and enable reliable detection of manipulation.
  • Selective reporting. Model providers declare their participation in specific evaluation streams prior to each round. Scores are computed centrally and automatically, precluding selective disclosure or omission of unfavorable results.
  • Accessibility and liveness. Validator onboarding is lightweight to support broad participation. The decentralized network of test contributors ensures that evaluations continue on a regular basis, even with fluctuating validator involvement.
  • Test quality. Each test set is peer-reviewed, with validator influence weighted by reputation. Validators submitting low-quality or biased content may have their collateral slashed and their contribution discounted, promoting sustained data quality.
4.4 Considerations towards Implementation
There are multiple considerations that remain to be considered in order to make such a system feasible in the future:
  • Scoring metrics. The implementation hinges on a fair scoring system that should be collusion-resistant and whose data quality scores can be fairly aggregated among heterogeneous metrics. We envision a reputation-based system to aggregate different validators’ scores - Provision of such a system is left up to future work.
  • Allowing data reuse. Our current draft reveals all test data after each round, which prevents the reuse of leaked test cases. However, it might be desirable to allow a certain degree of reuse in practice, as good tests are hard to come up with. The goal should be to only open some of the tests to the public at the end of each round, while maintaining fair scoring.
  • Scoring new models. For simplicity, we assumed that only models registered before the current round are scored. An ideal solution should score on a rolling basis, but exclude already revealed tests from evaluation.

### v2

@@ 3 Towards a New Paradigm: From Static Leaderboards to Proctored Exams
Note: A core challenge for achieving such a desired paradigm is temporal fairness.
  • If evaluation does not occur simultaneously on all evaluated models, then information can leak across models and periods (especially for models with the same creator but evaluated at different time) which induces contamination.
  • If test data are created after a model appears, then contributors can cherry pick items to favor or handicap that particular model, though possibly reputation systems, carefully designed mechanisms, and cross validation may help mitigate the risk.
As a result, the necessary conditions for the fairest score to arise include
  • (1) Tests are created before a model is released and remain fully secret until evaluation;
  • (2) All eligible models are evaluated at the same time on the same undisclosed items, and the items should be discarded for future evaluation once used for evaluation.
This ideal is extremely demanding for test creators to continuously deliver high-quality test cases, and it reduces direct comparison between models that do not appear in the same cohort.
The remainder of the paper instantiates some of these principles in a concrete, practical, and community-governed prototype, PeerBench, whose architecture and methodological safeguards are presented in Section 4, demonstrating the practicality of our position. PeerBench  prioritizes instant evaluation for new models over ultimate fairness, and reputation systems and slashing mechanisms are deployed for mitigating possible contamination risks.
4 PeerBench: A Live, Community Governed Benchmarking Platform
We introduce PeerBench [^2] Its prototype implementation is available at https://peerbench.ai., a platform that instantiates a proctored exam paradigm for model evaluation. The system combines a lightweight layer of community governance with a cryptographically verifiable workflow. The goal is sustained data quality for tests and a live evaluation that resists contamination and remains reproducible.
@@ 4.1 Actors
PeerBench supports multiple evaluation streams such as mathematics, code generation, and translation. Participants interact with a common coordination service and a live reservoir of tests.
Data contributors author private test suites together with executable scoring functions. They query registered models with their tests and record responses. Each contributor maintains a reputation that evolves through peer review of their tests.
Reviewers evaluate the quality of submitted tests. They produce ordinal ratings that determine test weights. Reviewer reputation is the Pearson correlation between their ratings and the final consensus quality of the same tests.
Model creators expose inference endpoints for their models and register for specific streams. Each model is evaluated exactly once on each test before that test retires and becomes public.
Coordination server authenticates uploads, manages the live reservoir, schedules peer review, updates reputations, cross validates scores against external benchmarks, and publishes public leaderboards. It stores all artifacts in a database and immutable object storage for audit.
End users are researchers, journalists, regulators, and practitioners who consult PeerBench for live leaderboards. They may apply trust thresholds that reflect their tolerance for uncertainty.
4.2 Three Leaderboards
The platform maintains three leaderboards that update continuously.
  1. 1.
    Data contributor leaderboard ranks contributors by cumulative test quality and verification bonuses
    $\text{ContributorScore}(c)\;=\;\sum_{i}\text{quality}\!\big(T^{(c)}_{i}\big)\;+\;\text{bonuses}.$ (1)
    This rule rewards both quality and steady contributions. Bonuses come from successful verification of track record of expertise (e.g. edu email, Google scholar profile, GitHub profile, etc.).
  2. 2.
    Reviewer leaderboard ranks reviewers by accuracy relative to consensus quality score
    $\text{ReviewerScore}(r)\;=\;\operatorname{Pearson}\!\left(\{q^{(i)}_{r}\},\{\overline{q}^{(i)}\}\right).$ (2)
    This rule rewards alignment with consensus across streams, and reviewers with higher reputations will be given priorities in reviewing new prompts, while malicious reviewers identified via consistently low reputation will be punished or removed from the platform.
  3. 3.
    Model leaderboard ranks models by a weighted average that respects test quality
    $\text{ModelScore}(m)\;=\;\frac{\sum_{i}w(T_{i})\,s^{(m)}_{i}}{\sum_{i}w(T_{i})}.$ (3)
    This rule takes data quality into the evaluation metric, striving to give a fair, robust, and contamination-free evaluation of the true abilities of different AI models.
4.3 Temporal Fairness Dilemma and Scheduling in PeerBench
As discussed in the previous section, a core challenge in PeerBench is temporal fairness. Models submitted at different times face different evaluation conditions as the test reservoir evolves, creating a fundamental tension between immediate evaluation and synchronized fairness.
Design Choice A: Immediate Scoring on Request
Under this paradigm, a model is scored immediately upon request. This approach maximizes responsiveness and enables continuous iteration, aligning with the widely adopted on-demand benchmarking paradigm. The system adapts naturally to fluctuating data contribution rates by maintaining only the most recent and highest-quality tests. The reputation system and weighted scoring help mitigate contamination effects through post-hoc identification of malicious actors. However, this approach still risks contamination from earlier interactions with similar items that may have leaked into training corpora, particularly for models from the same creator, and it cannot be fully and immediately addressed through reputation. Additionally, models evaluated at different times face different test sets, complicating direct comparison.
Design Choice B: Registration with Periodic Synchronized Evaluation
Models register for predetermined evaluation windows following a geometric progression (e.g., $2^{0},2^{1},2^{2},\ldots,2^{k}$ days). At each window’s end, all registered models are evaluated simultaneously on the same finalized secret test set. This guarantees the strongest form of fairness through secrecy and simultaneity, ensuring clear comparability within cohorts. However, it reduces inter-cohort comparability and increases operational burden on test creators. When data contribution fluctuates, platform liveness cannot be guaranteed as used tests are immediately made public and cannot be reused. Model creators must also wait for evaluation windows, potentially slowing development cycles.
Platform Stance
PeerBench adopts a hybrid approach, aiming to support both paradigms with a portion of data dedicated to immediate scoring and the remainder for synchronized evaluation. In the prototype, we prioritize immediate scoring for flexibility and timely feedback (i.e. Design Choice A). Synchronized windows produce gold standard cohort results, and cross-validation between both approaches provides additional confidence metrics. The coordination server records evaluation mode and cohort identifiers for every score, with public leaderboards displaying this metadata to enable informed interpretation of comparability. This hybrid strategy balances fairness, liveness, and practical utility while maintaining transparency about inherent trade-offs.
4.4 End-to-End Workflow of PeerBenchPrototype
Figure 1 in Appendix A presents the overall architecture and control flow. The design uses three visible lanes for contributors, the coordination server with the live reservoir, and reviewers. Models appear as external endpoints that receive queries and return responses through authenticated calls.
4.4.1 Setup Phase
Data contributors, reviewers, and model creators register with verifiable credentials such as institutional email. Each participant generates a public signing key. Contributors and reviewers stake collateral that can be slashed when misconduct is detected. Contributors receive initial reputation from verification bonuses. Reviewers begin with a neutral reputation.
4.4.2 Continuous Evaluation
The system maintains a live reservoir of size $k$. Only retired tests become public. A public test never influences evaluations of models that arrive after its publication.
T1. Test submission and commitment A contributor $c$ submits a test $T^{(c)}$ with an executable scoring function $F^{(c)}$. The system records a binding commitment $h=\operatorname{Com}(T^{(c)},F^{(c)})$ that prevents retroactive cherry picking and fixes the evaluation stream.
T2. Model evaluation The server schedules immediate queries to all currently registered models. The contributor stores audit logs and raw answers $A^{(c,m)}$ and computes objective scores $s^{(c,m)}$ using the provided scoring function. Each model is evaluated once on each live test. Evaluation must complete before retirement and publication.
T3. Review process The server assigns the test to reviewers at random with a requirement of at least three valid reviews. Reviewers assign integer quality scores $q\in\{-1,0,1,2\}$. Received reviews are visible to other reviewers for the same test. The final test quality $\overline{q}^{(i)}$ is the average of the collected scores weighted by the reputation of the corresponding reviewers.
T4. Weight calculation The final weight balances test quality and contributor reputation [^3] It’s an instance of the weight calculation formula, and can be subject to further discussion and change.
$w\!\big(T^{(c)}\big)\;=\;\max\!\left\{0,\;0.7*\,\text{quality}\!\big(T^{(c)}\big)+0.3*\,\min\!\big(2,\rho_{c}/100\big)\right\}.$ (4)
T5. Reservoir management The new test joins the live reservoir. The server retires tests at every step. Priority goes to tests with zero weight. If none exist the server retires the oldest test. Retirement triggers publication of the full artifact. This includes the test data, the score logs, the peer reviews, and the model responses. Retired tests never contribute to future evaluations.
T6. Reputation updates After each round, the reputation of all associated data contributors, reviewers, and models is updated according to Equations 1 2 3  respectively.
M1. New model integration When a new model is submitted to the platform, it is evaluated on the current reservoir for a preliminary initial score. However, the score may be affected by data contamination (e.g., exactly the same prompt may be leaked into the training corpora through earlier evaluation of another model of the same creator); and the score will eventually converge to a fair one once all tests that the model faced before its arrival are retired and published.
4.5 Security and Audit
Partial revelation for online peer review The system reveals small random portions of live tests to reviewers in a read-only and non-copyable format such as images. Reviewers verify consistency with the commitment $h$, and re-run scoring to confirm that published scores match the logs. Significant mismatch will trigger investigation and possible punishment. Finally, reviewers will assign a score based on the quality of the test (e.g., soundness, clarity, novelty, etc.).
Full publication After retirement the platform publishes the test $T^{(c)}_{i}$, the logs, and the model responses $A^{(c,m)}_{i}$. Anyone on the platform can check consistency with the hash commitments and flag any invalid prompts or possible misbehaviors. The central coordination server can take action on any misbehavior that successfully slips through the initial peer review but eventually gets caught by the community members after full publication.
Slashing mechanism and Economic model Participants whose reputation falls below a threshold are removed. In an ideal workflow, participants stake a certain amount of collateral before joining the platform, and receive rewards for positive contribution (reflected by the reputation scores on data contribution and reviewing). The platform slashes collateral for malicious tests or systematic deviation from consensus and uses the slashed collateral (and possible revenue if benchmarking on PeerBench  platform is made as a paid service to model creators) to reward honest contributors of the platform, creating a self-sustainable economic loop.
@@ 4.6 How Our Design Choices Address Common Issues
  • Data contamination and cherry picking. Validators pre-commit to test sets, which remain private until the round concludes and are never reused. This ensures that training on this data is infeasible if the validator behaves honestly. Any collusion or cherry-picking, such as tailoring questions to favored models, can be detected via cross-validator score discrepancies, discouraging misconduct through slashing and reputation penalties. This incentivizes honest behaviour.
  • Cheating on private data. A public source of randomness determines which queries are revealed, preventing validators from anticipating which items will be audited. Hash commitments to all related artifacts ensure verifiable consistency and enable reliable detection of manipulation.
  • Test quality. Each test receives multiple independent reviews. Data quality decides its weight in the final score of models. Low quality or biased content is discounted or removed, and continuously spamming the platform with low quality data may be subject to monetary punishment and slashing.
  • Accessibility and continuous operation. Registration is light for all roles which supports broad participation. Continuous operation maintains evaluation momentum when participation fluctuates. Reviewers require appropriate incentives for consistent participation, with monetary compensation or platform credits as plausible options that align with the reputation system.
@@ 6 Conclusion
Appendix A Workflow of PeerBench Prototype
Data contributorsCoordination server and live reservoirReviewers Submit private test $T^{(c)}$ with scoring function $F^{(c)}$ Publish hash commitment $h=\mathrm{Com}\!\big(T^{(c)},F^{(c)}\big)$ Authenticated queries to registered models Store logs and raw answers $A^{(c,m)}$ Compute objective scores $s^{(c,m)}$ Authenticate uploads and register test in stream Schedule immediate evaluation of all current models Model endpoints $M_{1},\ldots,M_{n}$ secure inference with rate limits and signed requests Materialize audit logs and cross check score computations Assign to reviewers size kkLive test reservoir $|R|>k$ Retirement prioritize low-quality zero weight tests. If none exists, retire the oldest test. Publish full artifacts on retirement. Update contributor, reviewer, and model leaderboards Cross validation with external benchmarks and reference sites Receive random assignment of tests at least three valid reviews required Assign quality $q\in\{-1,0,1,2\}$ reviewers can view received reviews Compute average quality of each test Compute the weight of each test into the final score Model leaderboard Contributor leaderboard Reviewer leaderboard Contributor reputation
$\rho_{c}=\sum\text{quality scores}+\text{bonuses}$
Reviewer reputation
$\rho_{r}=\text{average}\ \mathrm{Pearson}(\text{given ratings},\ \text{final quality})$
Continuous operation with rolling publication of retired tests and immutable audit artifacts
Figure 1: Overall architecture and continuous workflow in PeerBench. Blue denotes data contributors. Gray denotes the coordination service and the live reservoir. Green denotes reviewers. Orange denotes model endpoints. Dashed arrows show reputation feedback and cross validation links.